There is an economy with three firms, two of which make intermediate goods (Firm 1 produces an input for Firms 2, who produces an input for Firm 3), and one (Firm 3) which constructs a final good. Complete the following table.

                     Firm1   Firm 2   Firm 3    Totals

Sales                                                            ——–

Inputs             0          6000                        ——–

Labor           4,300     850       1,700

Rent             260         75                          650

Interest         200                         600

Profit                           750                      4,875

Value Added            2,000     5,500


Calculate GDP in this economy, and show that payments to factors of production, spending on final goods, and total value added are equal.



1
Expert's answer
2021-02-02T10:21:07-0500

Firm 1 sales = Firm 2 input = 6000.


Firm 1 profit = Firm 1 sales - sum of Firm 1 labor, rent and interest =6000 - (4300+260+200) = 1240.


Firm 1 sales = Value added for Firm 1 = 6000.


Total value added = sum of value added by each Firm = 6000+2000+5500 =13500.


Total value added = Sales by Firm 3 = 13500.


Total labor = Sum of each Firm’s labor =4300+850+1700 =6850.


Firm 3 rent = Total rent - (sum of Firm 1 rent and Firm 2 rent) = 650 - (260+75) =315.


Firm 3 profit = Total profit - (sum of Firm 1 profit and Firm 2 profit) =4875 - (1240+750) =2885.


Total interest = Total value added - ( sum of Total labor, Total rent and Total profit) = 13500 - (6850+650+4875) =1125.


Firm 2 interest = Total interest- (sum of Firm I interest and Firm 3 interest) = 1125 - (200+600) =325.


Firm 2 sales = Firm 3 sales - Firm 1 sales =13500 - 6000 =8000.


Firm 2 sales = Firm 3 input =8000.




The GDP in the economy = Total Value added = 6000+2000+5500 =13500.

Payment to factors of production = Total profit+ Total rent + Total interest + Total labor =6850+650+1125+4875 =13500.

Spending on final goods = Firm 3 sales =13500.


Total value added = Payment of factors of production = Spending on final goods =13500.